OpenStudy (anonymous):

A function is shown below: f(x) = x3 + 5x2 - x - 5 Part A: What are the factors of f(x)? Show your work. Part B: What are the zeros of f(x)? Show your work.

OpenStudy (gebooors):

Factors are 1, 5, -1, -5, also multiplier of terms. Remember sign!

OpenStudy (gebooors):

x3 + 5x2 - x -5 = 0 x2 ( x+ 5) - (x+ 5) = 0 Can you continue here?

OpenStudy (the_fizicx99):

Do you know how to factor this out?

OpenStudy (anonymous):

Not really /:

OpenStudy (anonymous):

wait so x3 + 5x2 - x -5 = 0 x2 ( x+ 5) - (x+ 5) = 0 is part b right @gebooors

OpenStudy (the_fizicx99):

Just because I need to learn better LaTeX I'll do it >.> 1. Split this 4 term polynomial into 2 groups \(\ x^3 + 5x^2 - x -5 \rightarrow (x^3 + 5x^2)\) and \(\ (-x - 5) \) Factor \(\ x^3 + 5x^2 \) Notice they both don't have coefficient gcf, however, they do have variable gcf, "x"! So if you factor a \(\ x^2\) from \(\ x^3 + 5x^2 \) you get \(\ x^2(x + 5) \) Now can you factor \(\ (-x - 5)\) ?

OpenStudy (anonymous):

so it wold be -x(-5) ?

OpenStudy (anonymous):

your response is in regards to part c right

OpenStudy (the_fizicx99):

Um no, there is no variable gcf or coefficient. So simply multiply it by -1. The reason why you're multiplying by -1 is because you want both terms to have a gcf. So if you use -1 You would get: \(\ -1(x+5) \) so \(\ -1(x+5) = -x - 5 \) Now both terms have a gcf, (x+5) ! \(\ x^2(x+5) -1(x+5) \) now this becomes \(\ (x+5)(x^2-1) \)

OpenStudy (the_fizicx99):

You have factored it out, now to find the zero's, solve for x :)

OpenStudy (the_fizicx99):

What's part C, no were doing it together, then \(\ \color{blue}{you're} \) going to answer them with what you have solved. Should be easy, I'll be guiding you

OpenStudy (anonymous):

if u say so

OpenStudy (the_fizicx99):

solve for "x"

OpenStudy (anonymous):

lol, good one

OpenStudy (the_fizicx99):

Can you try to solve?

OpenStudy (anonymous):

honestly, no, not even if a gun was held against me

OpenStudy (the_fizicx99):

o.o K, well... x + 5 = 0 and \(\ x^2 - 1 = 0 \) now?

OpenStudy (anonymous):

x2-1=0 +1 +1 x2=1

OpenStudy (the_fizicx99):

It's still x^2 = 1 continue, there's 1 more step (the inverse of the square is the square root)

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

well u lost me, suprise suprise

OpenStudy (anonymous):

wow, all of that to have the exact same answer except one less exponent.. YAY MATH

OpenStudy (the_fizicx99):

It won't always be like that...

OpenStudy (the_fizicx99):

x + 5 = 0

OpenStudy (anonymous):

x=5

OpenStudy (the_fizicx99):

Oh wait a minute! you could have easily applied the differences of squares! \(\ (x^2 - 1) = (x + 1)(x-1) \) So you have (x + 5)(x-1)(x+1) I'll just solve the zero's. x + 5 = 0 -5 -5 x = -5 x - 1 = 0 + 1 +1 x = 1 x + 1 = 0 -1 -1 x = -1 Overall: (-5,0) , (-1,0) and (1,0)

OpenStudy (anonymous):

ok so thats part b right

OpenStudy (the_fizicx99):

Those are your x intercepts (zero's) yes.

OpenStudy (the_fizicx99):

http://prntscr.com/36txhv

OpenStudy (the_fizicx99):

Well yeah, we just used one of the factors and solved for the zero's. Maybe @Zale101 can clarify any question. I have to get off. Night!

OpenStudy (anonymous):

thank you so much !!

OpenStudy (the_fizicx99):

Yw

OpenStudy (anonymous):

Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (anonymous):

@tHe_FiZiCx99 there was one mre part /:

OpenStudy (the_fizicx99):

Sorry, I had to get off :/